Calculate the mass of hydrogen formed when 25 grams of aluminum reacts with excess hydrochloric acid.

2Al + 6HCl → Al2Cl6 + 3H2
1
Expert's answer
2020-03-30T10:51:12-0400

According to the condition, hydrochloric acid is contained in excess in the reaction mixture. Thus, the amount of hydrogen substance should be calculated using the amount of aluminium substance:

  • "\u03bd(Al) = m\/M" = 25(g) / 26.98 (g/mol) = 0.93 mol
  • ν (H2) = "\u03bd (Al) \u2219 3\/2" = 0.93 (mol) ∙ 3/2 = 1.40 mol
  • m(H2) = "\u03bd\u2219M" = 1.40 ∙ 2.00 = 2.80 g

Answer: m(H2)=2.80 g